# Functions of bounded variation and Lipschitz algebras in metric measure spaces

### v1

@@ Acknowledgements
The first named author was supported by the Research Council of Finland grant 362898.
@@ 2. Preliminaries
Given any real number $a\in\mathbb{R}$, we denote by $a^{+}\coloneqq a\vee 0$ and $a^{-}\coloneqq a\wedge 0$ its positive part and its negative part, respectively. Given a set ${\rm X}$, we indicate by $\mathbbm{1}_{S}$ the characteristic function of a subset $S$ of ${\rm X}$, i.e. we define $\mathbbm{1}_{S}(x)\coloneqq 1$ for every $x\in S$ and $\mathbbm{1}_{S}(x)\coloneqq 0$ for every $x\in{\rm X}\setminus S$.
@@ 2.1. Metric/measure spaces
[…]
see e.g. [36, Lemma 2.1.27] (notice that, even if there the algebra $\mathscr{A}$ is assumed to be compatible, one can easily check that the simple proof of the mentioned Lemma does not require additional assumptions on $\mathscr{A}$). It can be readily checked that ${\rm LIP}_{\star}({\rm X})$ itself is a separating algebra.
@@
Fix $\delta>0$. Since $\mathscr{A}$ is dense in $L^{1}(\mathfrak{m})$, for any $i=1,\ldots,k$ we can find $(\psi^{k}_{i})_{k}\subseteq\mathscr{A}$ such that $\mathbbm{1}_{E_{i}}(x)=\lim_{k}\psi^{k}_{i}(x)$ for $\mathfrak{m}$-a.e. $x\in{\rm X}$. Thanks to [36, Corollary 2.1.24], we can also assume that $0\leq\psi^{k}_{i}\leq 1$. By Lemma 2.3, there is a polynomial $p$ such that
@@
2.2. $L^{p}$-Banach $L^{\infty}$-modules
@@
$\textsc{Int}_{\mathscr{M}}(\omega)(v)\coloneqq\int\omega(v)\,{\mathrm{d}}\mathfrak{m}\quad\text{ for every }\omega\in\mathscr{M}^{*}\text{ and }v\in\mathscr{M},$
@@
[…]
@@ 2.3. Lipschitz derivations
As $\mathscr{A}$ is dense in $L^{1}(\mathfrak{m})$ by (2.2), we have that ${\rm div}(b)$ is uniquely determined. In principle, ${\rm div}(b)$ may depend on the choice of the algebra $\mathscr{A}$. However, it is easy to check that if $\mathscr{A}\subseteq\tilde{\mathscr{A}}$ are unital separating subalgebras of ${\rm LIP}_{\star}({\rm X})$ and $b\in{\rm Der}^{\infty}_{\infty}({\rm X};\tilde{\mathscr{A}})$, then $b\in{\rm Der}^{\infty}_{\infty}({\rm X};\mathscr{A})$ and the two notions of divergence (in duality with $\mathscr{A}$ and $\tilde{\mathscr{A}}$, respectively) coincide. Thus, with a slight abuse of notation, we will write only ${\rm div}(b)$ without stressing the dependence on the chosen algebra, since it will be clear from the context. $\blacksquare$
@@ 3.1. The space ${\rm BV}_{\rm H}$
This is shown in Appendix A, reporting a detailed version of the arguments in the proofs of [4, Lemmata 5.2 and 5.4] showing that they do not make use of the completeness nor separability of the ambient metric space. Moreover, even though in [4, Eq. (5.1)] the approximating sequence in the definition of $|{\bf D}f|_{*}(U)$ is not required to consist of bounded functions, it still coincides with our definition thanks to a standard truncation argument.
@@
${\rm lip}_{a}(f_{n})\mathfrak{m}\rightharpoonup|{\bf D}f|_{*}\quad\text{ weakly.}$.
@@
It is highly non-trivial to prove that, assuming the ambient metric space is complete, every function in ${\rm BV}_{\rm H}({\rm X})$ can be approximated (on the whole ${\rm X}$) by a sequence of bounded Lipschitz functions. The validity of this claim follows from the results of [16], as we are going to see:
@@ 3.2. The space ${\rm H}^{1,1}$
$\|f\|_{{\rm H}^{1,1}({\rm X};\mathscr{A})}\coloneqq\|f\|_{L^{1}(\mathfrak{m})}+\||{\rm D}f|_{*,\mathscr{A}}\|_{L^{1}(\mathfrak{m})}\quad\text{ for every }f\in{\rm H}^{1,1}({\rm X};\mathscr{A}).$
@@ 3.3. The space ${\rm BV}_{\rm W}$
[…]
@@
$\|f\|_{{\rm BV}_{\rm W}({\rm X};\mathscr{A})}\coloneqq\|f\|_{L^{1}(\mathfrak{m})}+{\rm V}_{\mathscr{A}}(f)\quad\text{ for every }f\in{\rm BV}_{\rm W}({\rm X};\mathscr{A}).$
@@
Therefore, the measure ${\bf L}_{f}(b)$ is independent of the approximating sequence and $b(f_{n})\mathfrak{m}\rightharpoonup{\bf L}_{f}(b)$ holds for the original sequence $(f_{n})_{n}$. In particular, we have that ${\bf L}_{f}(b)$ is concentrated on ${\rm spt}(\mathfrak{m})$, proving (3.6c). If $b,\tilde{b}\in{\rm Der}^{\infty}_{\infty}({\rm X})$ and $\lambda\in\mathbb{R}$, then $(\lambda b+\tilde{b})(f_{n})\mathfrak{m}=\lambda\,b(f_{n})\mathfrak{m}+\tilde{b}(f_{n})\mathfrak{m}$ weakly converges both to ${\bf L}_{f}(\lambda b+\tilde{b})$ and $\lambda\,{\bf L}_{f}(b)+{\bf L}_{f}(\tilde{b})$, so that ${\bf L}_{f}$ is linear. Given $h,\tilde{h}\in{\rm LIP}_{\star}({\rm X})$, we have that $\int\tilde{h}\,{\mathrm{d}}{\bf L}_{f}(hb)=\lim_{n}\int\tilde{h}(hb)(f_{n})\,{\mathrm{d}}\mathfrak{m}=\lim_{n}\int(\tilde{h}h)b(f_{n})\,{\mathrm{d}}\mathfrak{m}=\int\tilde{h}h\,{\mathrm{d}}{\bf L}_{f}(b)=\int\tilde{h}\,{\mathrm{d}}(h\,{\bf L}_{f}(b))$, which yields (3.6b) thanks to (2.2). Moreover, for every $b\in{\rm Der}^{\infty}_{\infty}({\rm X})$ we can compute
@@ 4. Equivalence results
If $\mathscr{A}$ is a good algebra, then the approximation $(f_{n})_{n}$ of ${\sf d}_{x}\wedge r$ can be also chosen so that $0\leq f_{n}\leq r$ for every $n\in\mathbb{N}$. Indeed, letting $(f_{n})_{n}$ be as in Definition 4.1, we have that $u_{n}\coloneqq(f_{n}\wedge r)\vee 0\to{\sf d}_{x}\wedge r$ in $L^{1}(\mathfrak{m})$. Thanks to [36, Lemma 2.1.26], we can find a sequence $(\tilde{g}_{n})_{n}\subseteq\mathscr{A}$ such that $|\tilde{g}_{n}-u_{n}|\leq\frac{1}{n}$ and ${\rm lip}_{a}(\tilde{g}_{n})\leq{\rm lip}_{a}(u_{n})$ for all $n\in\mathbb{N}$. By the Dunford–Pettis theorem, we thus have (up to a non-relabelled subsequence) that ${\rm lip}_{a}(\tilde{g}_{n})\rightharpoonup\rho$ weakly in $L^{1}(\mathfrak{m})$, for some function $\rho\in L^{\infty}(\mathfrak{m})$ with $0\leq\rho\leq 1$. Finally, letting $g_{n}\coloneqq r\big{(}r+\frac{2}{n}\big{)}^{-1}\big{(}\tilde{g}_{n}+\frac{1}{n}\big{)}$ for every $n\in\mathbb{N}$, we obtain a sequence $(g_{n})_{n}\subseteq\mathscr{A}$ with $0\leq g_{n}\leq r$ such that $g_{n}\to{\sf d}_{x}\wedge r$ strongly in $L^{1}(\mathfrak{m})$ and ${\rm lip}_{a}(g_{n})\rightharpoonup\rho$ weakly in $L^{1}(\mathfrak{m})$, as desired. $\blacksquare$
@@
Notice that $\tilde{g}_{k}\in{\rm LIP}_{b}({\rm X})$ and ${\rm Lip}(\tilde{g}_{k})\leq L$. Since $\mathscr{A}$ is a good algebra, we can find $(f^{j}_{k})_{j}\subseteq\mathscr{A}$ such that $f^{j}_{k}\to\tilde{g}_{k}$ in $L^{1}(\mathfrak{m})$ and ${\rm lip}_{a}(f^{j}_{k})\rightharpoonup|{\rm D}\tilde{g}_{k}|_{*,\mathscr{A}}\leq L$ in $L^{1}(\mathfrak{m})$ as $j\to\infty$. By Remark 4.3, we can assume that $m\leq f^{j}_{k}\leq M$ for every $k,j\in\mathbb{N}$. Define
$g_{n}\coloneqq\tilde{g}_{1}\vee\ldots\vee\tilde{g}_{n}\in{\rm LIP}_{b}({\rm X})\quad\text{ for every }n\in\mathbb{N}.$
@@
    To prove it, take an approximating sequence $(f_{n})_{n}\subseteq\mathscr{A}$ for $f$ as in Proposition 4.4 and set $h_{n}\coloneqq({\rm lip}_{a}(f_{n})-{\rm Lip}(f))^{+}$. Letting $n\to\infty$ in $\bar{b}(f_{n})\leq|b|_{\mathscr{A}}{\rm lip}_{a}(f_{n})\leq|b|_{\mathscr{A}}({\rm Lip}(f)+h_{n})$, we get $\bar{b}(f)\leq{\rm Lip}(f)|b|_{\mathscr{A}}$. Doing the same for $-f$ instead, we obtain that $-\bar{b}(f)\leq{\rm Lip}(f)|b|_{\mathscr{A}}$, proving (4.5).
@@
We now define the pre-Cheeger $\mathscr{A}$-energy functional ~$\tilde{\mathcal{E}}_{\mathscr{A}}\colon L^{1}(\mathfrak{m})\to[0,+\infty]$ as
~$\tilde{\mathcal{E}}_{\mathscr{A}}(f)\coloneqq\left\{\begin{array}[]{ll}\int|{\rm D}f|_{*,\mathscr{A}}\,{\mathrm{d}}\mathfrak{m}\\ +\infty\end{array}\quad\begin{array}[]{ll}\text{ if }f\in\mathscr{A},\\ \text{ otherwise.}\end{array}\right.$
It can happen that ~$\tilde{\mathcal{E}}_{\mathscr{A}}|_{\mathscr{A}}\neq\mathcal{E}_{\mathscr{A}}|_{\mathscr{A}}$, but $\mathcal{E}_{\mathscr{A}}$ is the $L^{1}(\mathfrak{m})$-lower semicontinuous envelope of ~$\tilde{\mathcal{E}}_{\mathscr{A}}$ (thanks to (3.3) and to Remark 3.8). Notice also that the functional ~$\tilde{\mathcal{E}}_{\mathscr{A}}$ can be expressed as
~$\tilde{\mathcal{E}}_{\mathscr{A}}=\|\cdot\|_{L^{1}(T^{*}{\rm X};\mathscr{A})}\circ{\mathrm{d}}_{\mathscr{A}}.$ (4.6)
@@
${\sf n}_{\mathscr{A}}^{*}(V)=\left\{\begin{array}[]{ll}0\\ +\infty\end{array}\quad\begin{array}[]{ll}\text{ for every }V\in L^{1}(T^{*}{\rm X};\mathscr{A})^{\prime}\text{ with }\|V\|_{L^{1}(T^{*}{\rm X};\mathscr{A})^{\prime}}\leq 1,\\ \text{ for every }V\in L^{1}(T^{*}{\rm X};\mathscr{A})^{\prime}\text{ with }\|V\|_{L^{1}(T^{*}{\rm X};\mathscr{A})^{\prime}}>1.\end{array}\right.$ (4.7)
@@
Fix any $f\in{\rm BV}_{\rm W}({\rm X};\mathscr{A})$. Since $\mathcal{E}_{\mathscr{A}}$ is convex and $L^{1}(\mathfrak{m})$-lower semicontinuous, it coincides with the double Fenchel conjugate ~$\tilde{\mathcal{E}}_{\mathscr{A}}^{**}\coloneqq(\tilde{\mathcal{E}}_{\mathscr{A}}^{*})^{*}$ of ~$\tilde{\mathcal{E}}_{\mathscr{A}}$. Therefore, using (4.6), [10, Theorem 5.1], (4.7) and Lemma 4.7, we obtain that
$\begin{split}\mathcal{E}_{\mathscr{A}}(f)&=\tilde{\mathcal{E}}_{\mathscr{A}}^{**}(f)=\sup_{g\in L^{\infty}(\mathfrak{m})}\bigg{(}\int fg\,{\mathrm{d}}\mathfrak{m}-\tilde{\mathcal{E}}_{\mathscr{A}}^{*}(g)\bigg{)}=\sup_{g\in L^{\infty}(\mathfrak{m})}\bigg{(}\int fg\,{\mathrm{d}}\mathfrak{m}-({\sf n}_{\mathscr{A}}\circ{\mathrm{d}}_{\mathscr{A}})^{*}(g)\bigg{)}\\ &=\sup_{g\in L^{\infty}(\mathfrak{m})}\bigg{(}\int fg\,{\mathrm{d}}\mathfrak{m}-\inf_{\begin{subarray}{c}V\in D({\mathrm{d}}_{\mathscr{A}}^{*}):\\ {\mathrm{d}}_{\mathscr{A}}^{*}V=g\end{subarray}}{\sf n}_{\mathscr{A}}^{*}(V)\bigg{)}=\sup_{V\in D({\mathrm{d}}_{\mathscr{A}}^{*})}\bigg{(}\int f\,{\mathrm{d}}_{\mathscr{A}}^{*}V\,{\mathrm{d}}\mathfrak{m}-{\sf n}_{\mathscr{A}}^{*}(V)\bigg{)}\\ &\leq\sup_{\begin{subarray}{c}b\in{\rm Der}^{\infty}_{\infty}({\rm X};\mathscr{A}):\\ |b|_{\mathscr{A}}\leq 1\end{subarray}}\bigg{(}-\int f\,{\rm div}(b)\,{\mathrm{d}}\mathfrak{m}\bigg{)}=\sup_{\begin{subarray}{c}b\in{\rm Der}^{\infty}_{\infty}({\rm X};\mathscr{A}):\\ |b|_{\mathscr{A}}\leq 1\end{subarray}}{\bf L}_{f}(b)({\rm X})=|{\bf D}f|_{\mathscr{A}}({\rm X}).\end{split}$

### v2

@@ Acknowledgements
The first named author was supported by the Research Council of Finland grant 362898. We thank the reviewer for their useful comments, which have enhanced the quality of the manuscript.
@@ 2. Preliminaries
Given any real number $a\in\mathbb{R}$, we denote by $a^{+}\coloneqq a\vee 0$ and $a^{-}\coloneqq-(a\wedge 0)$ its positive part and its negative part, respectively. Given a set $\rm X$}, we indicate by $\mathbbm{1}_{S}$ the characteristic function of a subset $S$ of $\rm X$}, i.e. we define $\mathbbm{1}_{S}(x)\coloneqq 1$ for every $x\in S$ and $\mathbbm{1}_{S}(x)\coloneqq 0$ for every $x\in{\rm X}\setminus S$.
@@ 2.1. Metric/measure spaces
$\mathscr{A}\text{ is dense in }L^{p}(\mathfrak{m})\text{ for every }p\in[1,\infty),$ (2.2)
see e.g. [36, Lemma 2.1.27] (notice that, even if there the algebra $\mathscr{A}$ is assumed to be compatible, one can easily check that the simple proof of the mentioned Lemma does not require additional assumptions on $\mathscr{A}$). It can be readily checked that ${\rm LIP}_{\star}({\rm X})$ itself is a unital separating algebra. It is essentially because of the result in (2.2) that in the paper we will always assume that the algebras we are working with are unital and separating: both assumptions are indeed necessary to prove density results, which ultimately rely on the Stone–Weierstrass theorem. Furthermore, the fact that the algebras $\mathscr{A}$ that we consider contain all constant functions (as follows from the assumption that $\mathscr{A}$ is unital) will be implicitly used throughout the paper in several truncation arguments (for example, in the proof of Lemma 2.4). Indeed, when we apply [36, Corollary 2.1.24] to approximate a function of the form $f\wedge\lambda$, with $f\in\mathscr{A}$ and $\lambda\in\mathbb{R}$, with another function in $\mathscr{A}$ we rely on the fact that $\lambda\mathbbm{1}_{X}\in\mathscr{A}$. As we pointed out in the Introduction, allowing for constant functions in $\mathscr{A}$, we are forced to consider only finite reference measures $\mathfrak{m}$ (to make sure that all functions in $\mathscr{A}$ are integrable). It would be interesting, but outside the scopes of this paper and likely quite technical, to generalise some of the presented results to suitable (possibly non-unital) algebras and non-finite reference measures.
@@
Fix $\delta>0$. Since $\mathscr{A}$ is dense in $L^{1}(\mathfrak{m})$, for any $i=1,\ldots,n$ we can find $(\psi^{k}_{i})_{k}\subseteq\mathscr{A}$ such that $\mathbbm{1}_{E_{i}}(x)=\lim_{k}\psi^{k}_{i}(x)$ for $\mathfrak{m}$-a.e. $x\in{\rm X}$. Thanks to [36, Corollary 2.1.24], we can also assume that $0\leq\psi^{k}_{i}\leq 1$. By Lemma 2.3, there is a polynomial $p$ such that
@@
$2.$ $L^{p}$-Banach $L^{\infty}$-modules
@@
$\textsc{Int}_{\mathscr{M}}(\omega)(v)\coloneqq\int\omega(v)\,{\mathrm{d}}\mu\quad\text{ for every }\omega\in\mathscr{M}^{*}\text{ and }v\in\mathscr{M},$
@@
$\textsc{Int}_{\mathscr{M}}\text{ is an isometric isomorphism of Banach spaces.}$ (2.4)
@@ 2.3. Lipschitz derivations
[…]
@@ 3.1. The space ${\rm BV}_{\rm H}$
This is shown in Appendix A, explaining why the arguments in the proofs of [4, Lemmata 5.2 and 5.4] do not make use of the completeness nor separability of the ambient metric space. Moreover, even though in [4, Eq. (5.1)] the approximating sequence in the definition of $|{\bf D}f|_{*}(U)$ is not required to consist of bounded functions, it still coincides with our definition thanks to a standard truncation argument.
@@
${\rm lip}_{a}(f_{n})\mathfrak{m}\rightharpoonup|{\bf D}f|_{*}\quad\text{ weakly.}$
@@
It is highly non-trivial to prove that, assuming the ambient metric space is complete, every function in ${\rm BV}_{\rm H}({\rm X})$ can be approximated (on the whole $\rm X$}) as in Definition 3.1 by a sequence of bounded Lipschitz functions. The validity of this claim follows from the results of [16], as we are going to see:
@@ 3.2. The space ${\rm H}^{1,1}$
[…]
@@ 3.3. The space ${\rm BV}_{\rm W}$
${\bf L}_{f}(b)({\rm X})=-\int f\,{\rm div}(b)\,{\mathrm{d}}\mathfrak{m}$  $\quad\text{ for every }b\in{\rm Der}^{\infty}_{\infty}({\rm X};\mathscr{A}),$ (3.6a)
${\bf L}_{f}(hb)=h\,{\bf L}_{f}(b)$  $\quad\text{ for every }h\in\mathscr{A}\text{ and }b\in{\rm Der}^{\infty}_{\infty}({\rm X};\mathscr{A}),$ (3.6b)
$|{\bf L}_{f}(b)|({\rm X}\setminus{\rm spt}(\mathfrak{m}))=0$  $\quad\text{ for every }b\in{\rm Der}^{\infty}_{\infty}({\rm X};\mathscr{A}).$ (3.6c)
@@
[…]
@@
[…]
@@ 4. Equivalence results
If $\mathscr{A}$ is a good algebra, then the approximation $(f_{n})_{n}$ of ${\sf d}_{x}\wedge r$ can be also chosen so that $0\leq f_{n}\leq r$ for every $n\in\mathbb{N}$ and ${\rm lip}_{a}(f_{n})\to|{\rm D}({\sf d}_{x}\wedge r)|_{*,\mathscr{A}}$ strongly in $L^{1}(\mathfrak{m})$. To prove it, we first take a sequence $(g_{n})_{n}\subseteq\mathscr{A}$ such that $g_{n}\to{\sf d}_{x}\wedge r$ in $L^{1}(\mathfrak{m})$ and ${\rm lip}_{a}(g_{n})\leq G_{n}$ for some $(G_{n})_{n}$ strongly converging in $L^{1}(\mathfrak{m})$ to $|{\rm D}({\sf d}_{x}\wedge r)|_{*,\mathscr{A}}$, whose existence stems from Mazur’s lemma and the convexity of $f\mapsto{\rm lip}_{a}(f)$. Note that $u_{n}\coloneqq(g_{n}\wedge r)\vee 0\to{\sf d}_{x}\wedge r$ in $L^{1}(\mathfrak{m})$. Thanks to [36, Lemma 2.1.26], we can find a sequence $(\tilde{f}_{n})_{n}\subseteq\mathscr{A}$ such that $|\tilde{f}_{n}-u_{n}|\leq\frac{1}{n}$ and ${\rm lip}_{a}(\tilde{f}_{n})\leq{\rm lip}_{a}(u_{n})\leq{\rm lip}_{a}(g_{n})\leq G_{n}$ for all $n\in\mathbb{N}$. By the Dunford–Pettis theorem and the minimality of $|{\rm D}({\sf d}_{x}\wedge r)|_{*,\mathscr{A}}$, we thus have (up to a non-relabelled subsequence) that ${\rm lip}_{a}(\tilde{f}_{n})\rightharpoonup|{\rm D}({\sf d}_{x}\wedge r)|_{*,\mathscr{A}}$ weakly in $L^{1}(\mathfrak{m})$. In particular, testing against the constant function $\mathbbm{1}_{\rm X}\in L^{\infty}(\mathfrak{m})$ we deduce that $\int_{\rm X}{\rm lip}_{a}(\tilde{f}_{n})\,{\mathrm{d}}\mathfrak{m}\to\int_{\rm X}|{\rm D}({\sf d}_{x}\wedge r)|_{*,\mathscr{A}}\,{\mathrm{d}}\mathfrak{m}$, so that
$c_{n}\coloneqq\frac{\int_{\rm X}|{\rm D}({\sf d}_{x}\wedge r)|_{*,\mathscr{A}}\,{\mathrm{d}}\mathfrak{m}}{\int_{\rm X}{\rm lip}_{a}(\tilde{f}_{n})\,{\mathrm{d}}\mathfrak{m}}\to 1\quad\text{ as }n\to\infty.$
Therefore, the functions $\bar{f}_{n}\coloneqq c_{n}\tilde{f}_{n}\in\mathscr{A}$ satisfy $\bar{f}_{n}\to{\sf d}_{x}\wedge r$ strongly in $L^{1}(\mathfrak{m})$ as $n\to\infty$, along with $\int_{\rm X}{\rm lip}_{a}(\bar{f}_{n})\,{\mathrm{d}}\mathfrak{m}=\int_{\rm X}|{\rm D}({\sf d}_{x}\wedge r)|_{*,\mathscr{A}}\,{\mathrm{d}}\mathfrak{m}$ for every $n\in\mathbb{N}$. In particular, we have that
$\int_{\rm X}|{\rm lip}_{a}(\bar{f}_{n})-|{\rm D}({\sf d}_{x}\wedge r)|_{*,\mathscr{A}}|\,{\mathrm{d}}\mathfrak{m}$ $=2\int_{\rm X}({\rm lip}_{a}(\bar{f}_{n})-|{\rm D}({\sf d}_{x}\wedge r)|_{*,\mathscr{A}})^{+}\,{\mathrm{d}}\mathfrak{m}$
$\leq\int_{\rm X}(c_{n}G_{n}-|{\rm D}({\sf d}_{x}\wedge r)|_{*,\mathscr{A}})^{+}\,{\mathrm{d}}\mathfrak{m}\to 0.$
Finally, letting $f_{n}\coloneqq c_{n}^{-1}r\big(r+\frac{2}{n}\big)^{-1}\big(\bar{f}_{n}+\frac{c_{n}}{n}\big)$ for all $n\in\mathbb{N}$, we obtain the desired sequence. $\blacksquare$
@@
Notice that $\tilde{g}_{k}\in{\rm LIP}_{b}({\rm X})$ and ${\rm Lip}(\tilde{g}_{k})\leq L$. Since $\mathscr{A}$ is a good algebra, by Remark 4.3, we can find $(f^{j}_{k})_{j}\subseteq\mathscr{A}$ such that $f^{j}_{k}\to\tilde{g}_{k}$, ${\rm lip}_{a}(f^{j}_{k})\to|{\rm D}\tilde{g}_{k}|_{*,\mathscr{A}}\leq L$ in $L^{1}(\mathfrak{m})$ as $j\to\infty$, and $m\leq f^{j}_{k}\leq M$ for every $k,j\in\mathbb{N}$. Define
[…]
@@
The following two results contain the first steps to prove the equivalence of definitions of BV spaces in Theorem 4.10, in particular the inclusion ${\rm BV}_{\rm W}({\rm X};{\rm LIP}_{b}({\rm X}))\subseteq{\rm BV}_{\rm W}({\rm X};\mathscr{A})$ contained in Corollary 4.6, which is an immediate consequence of the next Theorem 4.5.
@@
    To prove it, take an approximating sequence $(f_{n})_{n}\subseteq\mathscr{A}$ for $f$ as in Proposition 4.4 and set $h_{n}\coloneqq({\rm lip}_{a}(f_{n})-{\rm Lip}(f))^{+}$. Letting $n\to\infty$ in $\bar{b}(f_{n})\leq|b|_{\mathscr{A}}{\rm lip}_{a}(f_{n})\leq|b|_{\mathscr{A}}({\rm Lip}(f)+h_{n})$, we get $\bar{b}(f)\leq{\rm Lip}(f)|b|_{\mathscr{A}}$. Doing the same for f-f instead, we obtain that $\bar{b}(f)\leq{\rm Lip}(f)|b|_{\mathscr{A}}$, proving (4.5).
@@
We now define the pre-Cheeger $\mathscr{A}$-energy functional $\tilde{\mathcal{E}}_{\mathscr{A}}\colon L^{1}(\mathfrak{m})\to[0,+\infty]$ as
$\tilde{\mathcal{E}}_{\mathscr{A}}(f)\coloneqq\left\{\begin{array}[]{ll}\int|{\rm D}f|_{*,\mathscr{A}}\,{\mathrm{d}}\mathfrak{m}\\ +\infty\end{array}\quad\begin{array}[]{ll}\text{ if }f\in\mathscr{A},\\ \text{ otherwise.}\end{array}\right$.
It can happen that $\tilde{\mathcal{E}}_{\mathscr{A}}|_{\mathscr{A}}\neq\mathcal{E}_{\mathscr{A}}|_{\mathscr{A}}$, but $\mathcal{E}_{\mathscr{A}}$ is the $L^{1}(\mathfrak{m})$-lower semicontinuous envelope of $\tilde{\mathcal{E}}_{\mathscr{A}}$ (thanks to (3.3) and to Remark 3.8). Notice also that the functional $\tilde{\mathcal{E}}_{\mathscr{A}}$ can be expressed as
$\tilde{\mathcal{E}}_{\mathscr{A}}=\|\cdot\|_{L^{1}(T^{*}{\rm X};\mathscr{A})}\circ{\mathrm{d}}_{\mathscr{A}}.$ (4.6)
@@
[…]
@@
Fix any $f\in{\rm BV}_{\rm W}({\rm X};\mathscr{A})$. Since $\mathcal{E}_{\mathscr{A}}$ is convex and $L^{1}(\mathfrak{m})$-lower semicontinuous, it coincides with the double Fenchel conjugate $\tilde{\mathcal{E}}_{\mathscr{A}}^{**}\coloneqq(\tilde{\mathcal{E}}_{\mathscr{A}}^{*})^{*}$ of $\tilde{\mathcal{E}}_{\mathscr{A}}$. Therefore, using (4.6), [10, Theorem 5.1], (4.7) and Lemma 4.7, we obtain that
$\begin{split}\mathcal{E}_{\mathscr{A}}(f)&=\tilde{\mathcal{E}}_{\mathscr{A}}^{**}(f)=\sup_{g\in L^{\infty}(\mathfrak{m})}\bigg(\int fg\,{\mathrm{d}}\mathfrak{m}-\tilde{\mathcal{E}}_{\mathscr{A}}^{*}(g)\bigg)=\sup_{g\in L^{\infty}(\mathfrak{m})}\bigg(\int fg\,{\mathrm{d}}\mathfrak{m}-({\sf n}_{\mathscr{A}}\circ{\mathrm{d}}_{\mathscr{A}})^{*}(g)\bigg)\\ &=\sup_{g\in L^{\infty}(\mathfrak{m})}\bigg(\int fg\,{\mathrm{d}}\mathfrak{m}-\inf_{\begin{subarray}{c}V\in D({\mathrm{d}}_{\mathscr{A}}^{*}):\\ {\mathrm{d}}_{\mathscr{A}}^{*}V=g\end{subarray}}{\sf n}_{\mathscr{A}}^{*}(V)\bigg)=\sup_{V\in D({\mathrm{d}}_{\mathscr{A}}^{*})}\bigg(\int f\,{\mathrm{d}}_{\mathscr{A}}^{*}V\,{\mathrm{d}}\mathfrak{m}-{\sf n}_{\mathscr{A}}^{*}(V)\bigg)\\ &=\sup_{\begin{subarray}{c}v\in L^{\infty}(T{\rm X};\mathscr{A}):\\ \textsc{Int}_{\mathscr{A}}(v)\in D({\mathrm{d}}_{\mathscr{A}}^{*}),\,\|v\|_{L^{\infty}(T{\rm X};\mathscr{A})}\leq 1\end{subarray}}\bigg(\int f\,{\mathrm{d}}_{\mathscr{A}}^{*}(\textsc{Int}_{\mathscr{A}}(v))\,{\mathrm{d}}\mathfrak{m}\bigg)\\ &=\sup_{\begin{subarray}{c}v\in L^{\infty}(T{\rm X};\mathscr{A}):\\ \textsc{Int}_{\mathscr{A}}(v)\in D({\mathrm{d}}_{\mathscr{A}}^{*}),\,\|v\|_{L^{\infty}(T{\rm X};\mathscr{A})}\leq 1\end{subarray}}\bigg(-\int f\,{\rm div}(b_{v})\,{\mathrm{d}}\mathfrak{m}\bigg)\\ &\leq\sup_{\begin{subarray}{c}b\in{\rm Der}^{\infty}_{\infty}({\rm X};\mathscr{A}):\\ |b|_{\mathscr{A}}\leq 1\end{subarray}}\bigg(-\int f\,{\rm div}(b)\,{\mathrm{d}}\mathfrak{m}\bigg)=\sup_{\begin{subarray}{c}b\in{\rm Der}^{\infty}_{\infty}({\rm X};\mathscr{A}):\\ |b|_{\mathscr{A}}\leq 1\end{subarray}}{\bf L}_{f}(b)({\rm X})=|{\bf D}f|_{\mathscr{A}}({\rm X}).\end{split}$